OpenStudy (amonoconnor):

I need helping evaluating the following integral: Integ. ((5x + 4)^(1/2))dx, from x=0 to 1 I am getting something different from what WolframAlpha says it should be equal to be, but I don't know where I'm going wrong. Would anyone be willing to solve this one themselves, and cross their work with mine? I got: 19/5 WolframAlpha: 38/15 Any and all help is greatly appreciated! :)

OpenStudy (anonymous):

To be clear, this is what you want? \[\int\limits_{0}^{1}\sqrt{5x+4}dx\]

OpenStudy (amonoconnor):

Yes! :)

OpenStudy (anonymous):

For the definite integral I get 4*(5*x + 4)**1.5/30 using formal methods (sympy, a Python library). Setting this expression to i(x) and evaluating i(1.)-i(0) yields 2.533333333333327, which is the W-A result.

OpenStudy (amonoconnor):

I used u-substitution, setting u = (5x+4). Can you help me verify the accuracy of my steps with that method? :)

OpenStudy (amonoconnor):

u = (5x + 4) du = (5x^(1-1) + (0)) dx = (5(x^0)) dx = (5)dx dx = (1/5)du So.... 1/5(integ.) u^1/2 * du is what I get for the integral, after u-sub, and before taking the anti-deriviative and plugging the numerical value of u back in.

OpenStudy (amonoconnor):

At that point This is what I did... 1/5(u^(.5 + 1)) = 1/5(u^(3/2)) = 1/5(5x + 4)^(3/2) Plugging in the limits... = [1/5(5(1) + 4)^(3/2)] - [1/5(5(0) + 4)^(3/2)] = [1/5(9)^(3/2)] - [1/5(4)^(3/2)] = [1/5(27)] - [1/5(8)] = [27/5] - [8/5] = 19/5 = 3.8 Where am I going wrong?? :/

OpenStudy (anonymous):

Please see attached. To be evaluated from upper limit of 9 to lower limit of 4.

OpenStudy (anonymous):

Quick read : have you rewritten the limits of integration in terms of u?

OpenStudy (amonoconnor):

...the limits?

OpenStudy (anonymous):

x goes from 0 to 1, so what will u do?

OpenStudy (anonymous):

Is that it? I will be away for a while.

OpenStudy (amonoconnor):

Well, with u-sub, doesn't one plug in the values of the limits to the "anti-derived" integral, with u plugged back in? This done by subtracting the quantity found using the lower from the quantity found using the upper?

OpenStudy (anonymous):

We're probably talking about different approaches. Once you have transformed from x to u you need not return to x, as long as you transform all parts of the original integral in terms of u. In this case, dx becomes du/5; the integrand becomes simply the square root of u; the limits of integration become 4 and 9 (rather than 0 and 1).

ganeshie8 (ganeshie8):

At that point This is what I did... 1/5(u^(.5 + 1)) = 1/5(u^(3/2)) ****************************************** = 1/5(5x + 4)^(3/2) Plugging in the limits... = [1/5(5(1) + 4)^(3/2)] - [1/5(5(0) + 4)^(3/2)] = [1/5(9)^(3/2)] - [1/5(4)^(3/2)] = [1/5(27)] - [1/5(8)] = [27/5] - [8/5] = 19/5 = 3.8 Where am I going wrong?? :/

ganeshie8 (ganeshie8):

mistake in that line, rest everything looks perfect

ganeshie8 (ganeshie8):

you forgot to divide (.5 + 1)

OpenStudy (amonoconnor):

I don't know what BillBell was talking about...

ganeshie8 (ganeshie8):

\(\large \int u^{\frac{1}{2}} du = \frac{u^{\frac{1}{2}+1}}{\frac{1}{2}+1}\) right ?

OpenStudy (amonoconnor):

Ah... Right *whoops!

OpenStudy (amonoconnor):

If I could double-fan you, I would!

ganeshie8 (ganeshie8):

so that 3/2 in the denominator is missing all the way to the end... to fix ur answer, simply multiply ur final answer by 2/3

OpenStudy (amonoconnor):

Thank you soo soo much !

ganeshie8 (ganeshie8):

np :) u will be seeing @BillBell 's method shortly i guess whenever ur professor takes time to explain it lol...

OpenStudy (amonoconnor):

Thank you again!